The spot exchange rate is S(¥120/$1). The U.S. interest rate is 4 percent APR. The Japanese interest rate is 2 percent APR. What is the no-arbitrage 90-day forward rate?

Multiple Choice

  • ¥119.4059/$1

  • ¥122.3529/$1

  • ¥120.5970/$1

  • None of the options

  • ¥117.6923/$1
Answer #1

(1+ix) = F(x/y) / S(x/y) * (1+iY)

1.02 = F(x/y) / 120 * 1.04

F(x/y) = (1.02*120) / 1.04

F(x/y) = 117.6923

Option 5 is correct